A small group of four is chosen randomly from a team of six. How many unique groups of four could be made?
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

There are 6*5*4*3 = 30*12 = 360 ways to pick 4 people where order matters.

But order doesn't matter, so we divide by 4! = 4*3*2*1 = 24 to get 360/24 = 15

Therefore, there are 15 different groups possible.

Answer by stanbon(75887)

A small group of four is chosen randomly from a team of six. How many unique groups of four could be made?
----------
Ans: 6C4 = 6!/[(6-4)!*4!] = (6*5)/(1*2) = 15 unique groups
=================
Cheers,
Stan H.
-----------------